# BatteryLife: A Comprehensive Dataset and Benchmark for Battery Life Prediction

### v6

@@ B.1. Further Details on Hyperparameters
For each model, we tried more than 20 different sets of hyperparameters and selected the hyperparameters that led to the lowest MAPE on the validation sets. The search ranges of hyperparameters are described as follows: batch size was chosen from [16, 32, 64, 128], learning rate was selected from [$5\times 10^{-4}$, $5\times 10^{-3}$, $1\times 10^{-3}$, dropout rate was chosen from [0, 0.05, 0.1], and embedding dimension was selected from [32, 64, 128, 256]. Additionally, for CyclePatch models, both the number of intra-cycle encoder layers and the number of inter-cycle encoder layers were tuned from 0 to 12. The number of CNN layers is tuned from 1 to 8.
@@ B.2. Implementation Details of Baselines
Regarding the CNN, we first reshape $\overline{\textbf{X}}_{1:100}$ into the shape of $\textbf{S}\in\mathbb{R}^{3\times 100\times 300}$, where 3 indicates the three variables (voltage, current and capacity), 100 is the number of maximum input cycles, 300 denotes the resampled length. Then, a CNN is utilized to model the S. In each layer, a Conv2d is first used, followed by ReLU activation. After the activation, average pooling is used to extract useful information. After a preset number of layers, a linear projection is used to produce the prediction.
As for the DLinear, PatchTST, Autoformer, iTransformer, and MICN, we use the implementations in https://github.com/thuml/Time-Series-Library. More specifically, we slightly revise the classification implementations to make models usable for BLP.

### v7

@@ B.1. Further Details on Hyperparameters
For each model, we tried more than 20 different sets of hyperparameters and selected the hyperparameters that led to the lowest MAPE on the validation sets. The search ranges of hyperparameters are described as follows: batch size was chosen from [16, 32, 64, 128], learning rate was selected from $5\times 10^{-4}$, $5\times 10^{-3}$, $1\times 10^{-3}$, dropout rate was chosen from [0, 0.05, 0.1], and embedding dimension was selected from [32, 64, 128, 256]. Additionally, for CyclePatch models, both the number of intra-cycle encoder layers and the number of inter-cycle encoder layers were tuned from 0 to 12. The number of CNN layers is tuned from 1 to 8.
@@ B.2. Implementation Details of Baselines
Regarding the CNN, we first reshape $\overline{\textbf{X}}_{1:100}$ into the shape of $\textbf{S}\in\mathbb{R}^{3,100,300}$, where 3 indicates the three variables (voltage, current and capacity), 100 is the number of maximum input cycles, 300 denotes the resampled length. Then, a CNN is utilized to model the S. In each layer, a Conv2d is first used, followed by ReLU activation. After the activation, average pooling is used to extract useful information. After a preset number of layers, a linear projection is used to produce the prediction.
As for the DLinear, PatchTST, Autoformer, iTransformer, and MICN, we use the implementation in https://github.com/thuml/Time-Series-Library. Specifically, we slightly revise the classification implementations to make models usable for BLP.